OpenStudy (anonymous):

Keisha's parents want to save twenty thousand dollars in her college savings account over the next fifteen years. They have eight thousand dollars to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth.

OpenStudy (anonymous):

a)2.5 b)2.7 c)10% d)12%

OpenStudy (anonymous):

Compound Annual Growth Rate (CAGR) is one of those terms best defined by example. Suppose you invested $10,000 in a portfolio on Jan 1, 2005. Let's say by Jan 1, 2006, your portfolio had grown to $13,000, then $14,000 by 2007, and finally ended up at $19,500 by 2008. Your CAGR would be the ratio of your ending value to beginning value ($19,500 / $10,000 = 1.95) raised to the power of 1/3 (since 1/# of years = 1/3), then subtracting 1 from the resulting number: 1.95 raised to 1/3 power = 1.2493. (This could be written as 1.95^0.3333). 1.2493 - 1 = 0.2493 Another way of writing 0.2493 is 24.93%. Thus, your CAGR for your three-year investment is equal to 24.93%, representing the smoothed annualized gain you earned over your investment time horizon.